1. A car traveling 36km/hr accelerates uniformly at 2m/s.calculate the velocity after 5s
1
Expert's answer
2021-06-30T10:14:39-0400

Gives

t=5sec

acceleration (a)=2m/sec


"u=36km/h=\frac{36\times10^{3}}{3600}=10m/sec"

V=?

We know that


"v=u+at=10+2\times5=20m/sec"

"v=20m\/sec"